A man age is 4 times the total age of of 3 sons .after 8 yrs ,his age would be twice the total age of his 3 sons .if the total present of 3 sons is x yrs . find the age of man
Answer by solve_for_x(190)

If the father's age is four times the current sum of the children's ages, x, then
the father's age can be represented by 4x.

In 8 years, the father's age will be 4x + 8.

The sum of the children's ages will be x + 24, since each of the three children will be 8 years older.

Since the father's age will be twice the sum of the children's ages, you can write:

4x + 8 = 2(x + 24)

4x + 8 = 2x + 48

4x - 2x + 8 = 48

2x + 8 = 48

2x = 48 - 8

2x = 40

x = 20

The current age of the man is 4x = 4(20) = 80 years old.